# SegmentDreamer: Towards High-fidelity Text-to-3D Synthesis with Segmented Consistency Trajectory Distillation

### v1

@@ Abstract
Figure 1: Examples of 3D assets generated by SegmentDreamer. Our framework addresses the improper conditional guidance issues of recent Consistency Distillation (CD)-based methods and theoretically provides a significantly tighter upper bound on distillation error, enabling high-fidelity text-to-3D generation within a reasonable timeframe ($\sim$32 minutes with classifier-free guidance [9] and $\sim$38 minutes combined with Perp-Neg [1]) on a single A100 GPU through 3D Gaussian Splatting (3DGS). Project Page: https://zjhJOJO.github.io
@@ 1 Introduction
Text-to-3D generation aims to create 3D assets that accurately align with text descriptions. It plays a vital role in various domains such as virtual reality, 3D gaming, and film production. One compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34] through Score Distillation Sampling (SDS) [31, 46], which leverages pre-trained text-to-image DMs as the 3D priors to guide the optimization of differentiable 3D representations, such as Neural Radiance Fields (NeRF) [29] and 3D Gaussian Splatting (3DGS) [17]. Subsequent research focus on improving the performance of SDS in diversity [47, 11, 43], visual quality [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16], and multi-view consistency [5, 10, 37, 35, 12].
@@
Figure 2: Visual example of existing CD-based methods: CDS [48], GCS [22], and our SegmentDreamer with a Classifier-Free Guidance (CFG) scale of 7.5. As shown, improper conditional guidance leads to suboptimal results. (a) Ineffective guidance produces semantically inconsistent details, while (b) excessive guidance results in overexposure and artifacts. Although (c) Brightness-Equalized Generation (BEG) [22] can mitigate overexposure, artifacts persist. In contrast, our SegmentDreamer enhances conditional generation, producing high-fidelity 3D assets.
@@
Recent methods [48, 22] have improved SDS and its variants by incorporating Consistency Distillation (CD) [40, 18] into their distillation loss. However, due to an imbalance between self-consistency and cross-consistency in their losses, these methods inherently suffer from improper conditional guidance when combined with Classifier-Free Guidance (CFG) for generation, ultimately leading to suboptimal outcomes (see Figure 2, a detailed explanation is provided in Section 4.1). Additionally, the distillation error upper bound of these methods remains large, limiting their generation quality.
@@
To address the aforementioned limitations, we present a novel framework – SegmentDreamer, which aims to fully unleash the potential of consistency models for high-fidelity text-to-3D generation. Specifically, by analyzing existing CD-based methods, we draw inspiration from segmented consistency models [33, 44] to reformulate SDS through the proposed Segmented Consistency Trajectory Distillation (SCTD), which effectively mitigates the imbalance between self-consistency and cross-consistency by explicitly defining their relationship. Moreover, unlike CD-based methods that directly enforce consistency across the entire Probability Flow Ordinary Differential Equation (PF-ODE) trajectory, SCTD partitions the PF-ODE trajectory into multiple sub-trajectories, ensuring consistency within each segment. This approach theoretically provides a significantly tighter upper bound on distillation error. In addition, we propose a distillation pipeline for a more swift and stable generation. We also introduce 3DGS as our 3D representation. Extensive experimental results demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality. Moreover, the proposed SCTD can be seamlessly applied to various 3D generative tasks (e.g., 3D avatar generation, and 3D portrait generation), facilitating the creation of high-fidelity 3D assets.
@@ 2.1 Text-to-3D Generation
Creating a 3D object from a given text description is a long-standing challenge in the fields of computer vision and graphics. To achieve this goal, a series of end-to-end text-to-3D pipelines have been proposed [8, 7, 30, 13, 21, 49], but these methods require large-scale labeled 3D datasets for training. To reduce the reliance degree of 3D training data, a compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34]. A groundbreaking contribution in this domain is DreamFusion [31], which introduced Score Distillation Sampling (SDS) loss to optimize parametric 3D representations. Follow-up works focus on refining and extending SDS in various aspects. For example, several studies [47, 11, 43] improve SDS in diversity through prompting modifications, while [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16] enhance the fidelity of the generated 3D objects. Other advancements aim to achieve multi-view consistency [5, 10, 37, 35, 12] and faster generation [42, 52]. Recently, Consistency Distillation (CD) [40, 18] has received widespread attentions as it enables high-quality generation through minimal sampling steps. Inspired by CD, Wu et al. [48] and Li et al. [22] incorporate CD into their distillation loss, which addresses some limitations brought by SDS. However, these CD-based methods inherently suffer from improper conditional guidance, ultimately leading to sub-optimal results. In this work, we aim to unleash the potential of CD to generate high-fidelity 3D assets, pushing the boundaries of text-to-3D synthesis.
@@ 2.2 Differentiable 3D Representations
Neural Radiance Field (NeRF) [29] aims to construct implicit 3D representations using a multi-layer perceptron (MLP), leveraging standard volumetric rendering [14] and alpha compositing techniques [32]. Subsequent research has explored adapting NeRF to various applications, Subsequent research has explored adapting NeRF to various applications, including sparse-view reconstruction [45], anti-aliasing [2], and medical image super-resolution [4].
@@ 3 Preliminaries
Diffusion Model (DM) [39, 38, 15] consists of a diffusion process and a denoising process. The diffusion process can be modeled as the solution $\hat{\text{o}}$ stochastic differential equation (SDE) with the transition kernel $\mathbb{P}(\mathbf{z}_{t}|\mathbf{z}_{0}):=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{z}_{0},\sigma_{t}\mathbf{I})$. By reversing the diffusion process, $\mathbf{z}_{0}$ can be estimated by solving a Probability Flow Ordinary Differentiable Equation (PF-ODE), which is defined by $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})]\mathrm{d}t$. A neural network ${\boldsymbol{\epsilon_{\phi}}}(\mathbf{z},t)$ can be trained to approximate the scaled score function $-\sigma_{t}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})$, offering an empirical formulation of the PF-ODE: $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}+\frac{g(t)^{2}}{2\sigma_{t}}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t)]\mathrm{d}t.$
Consistency Model [40, 27] aims to achieve few-step inference by distilling a pre-trained DM $\boldsymbol{\phi}$ into a student model $\boldsymbol{\phi}^{\prime}$. Given a PF-ODE trajectory $\{\mathbf{z}_{t}\}_{t\in[0,T]}$ and a condition $\mathbf{y}$, a consistency model $\boldsymbol{f}_{\boldsymbol{\phi}^{\prime}}$ builds the mapping $\boldsymbol{f}_{\boldsymbol{\phi}^{\prime}}(\mathbf{z}_{t},t,\mathbf{y})=\mathbf{z}_{0}$ between any points along this trajectory and the endpoint $\mathbf{z}_{0}$ by minimizing the following Consistency Distillation (CD) loss:
$\mathcal{L}_{\text{CD}}=\mathbb{E}[\lambda(t)||\boldsymbol{f}_{\boldsymbol{\phi}^{\prime}}(\mathbf{z}_{t},t),\boldsymbol{f_{\phi_{s}^{-}}}(\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s)||_{2}^{2}],t>s,$ (1)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ for $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\boldsymbol{\phi_{s}^{-}}$ is updated with the exponential moving average [40, 27]. $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ can be obtained by:
@@
The consistency function $\boldsymbol{f_{\phi_{s}}}$ can be parameterized in various ways [40, 27], and the solver $\boldsymbol{\Phi}(\cdot,\cdot,\cdot,\cdot)$ is used to calculate the point at $s$, which can be implemented using DDIM [38] or DPM-solver [26].
Score Distillation Sampling (SDS) Loss. Given a 3D representation $\boldsymbol{\theta}$ that can render a view image $\mathbf{z}_{0}:=\boldsymbol{g}(\mathbf{c},\boldsymbol{\theta})$ with camera pose $\mathbf{c}$, SDS distills a 2D DM $\boldsymbol{\phi}$ to generate a 3D object. by minimizing the following SDS loss w.r.t. $\boldsymbol{\theta}$:
@@
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$. Poole et al. [31] empirically omit the U-Net Jacobian term $\frac{\partial\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\mathbf{z}_{t}}$ in gradient calculation for speed-up, leading to low-fidelity results.
@@ 4 Our Method
In this section, we first review the mechanism of existing Consistency Distillation (CD)-based methods [48, 22] and reveal that they inherently suffer from improper conditional guidance when combined with Classifier-free Guidance (CFG). Subsequently, we address these limitations by reformulating SDS through the lens of the proposed Segmented Consistency Trajectory Distillation (SCTD). Moreover, building on this insight, we propose the efficient SCTD sampling for high-fidelity text-to-3D generation. We further tailor a distillation pipeline for the proposed SCTD sampling to achieve more swift and stable optimization. Finally, we provide a theoretical proof for our distillation error upper bound. The overall framework of our SegmentDreamer is depicted in Fig. 3, with the algorithm detailed in Algorithm 1.
@@ 4.1 Reviewing CD-based Losses
The consistency function w.r.t. 3D representation $\boldsymbol{\theta}$ is parameterized by $\boldsymbol{F_{\theta}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\mathbf{z}_{t}-\sigma_{t}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\alpha_{t}}$ [48, 22]. Before proceeding, we clarify two key concepts: self-consistency and cross-consistency. The former ensures that all points on the same ODE trajectory consistently map to their original points, as in consistency models [27, 40, 18]. The latter enforces an alignment between unconditional and conditional ODE trajectories [22].
@@
where $s<t$, $c(t)=\omega(t)(\frac{\alpha_{t}}{\sigma_{t}})^{2}$, and $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ is estimated using single-step DDIM sampling. Eq. 4 reveals that CDS prioritizes self-consistency while omitting cross-consistency in its loss function, which can be interpreted as the absence of the “classifier score”. As reported in CSD [53], this missing component, crucial for conditional generation, hinders CDS from producing semantically consistent content (see Fig. 5 for details). Moreover, the reliance on $\boldsymbol{F}_{\theta}$ may introduce a significant distillation error upper bound $\mathcal{O}(\Delta_{t})T$, where $\Delta_{t}=\max{|t-s|}$. This issue is particularly pronounced in text-to-3D settings due to the typically large $|t-s|$.
@@
where $e$$<$$s$$<$$t$. Moreover, GCS employs a conditional guidance loss $\mathcal{L}_{\text{CG}}$ and a pixel-wise constraint $\mathcal{L}_{\text{CP}}$ to enforce cross-consistency (see Supp. A for details).
However, GCS loss suffers from improper conditional guidance due to two key issues: (1) $\boldsymbol{G_{\theta}}$ used to enforce cross-consistency is inherently flawed due to the absence of a target time-step in $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, as detailed in App. B. (2) GCS employs $\boldsymbol{F}_{\boldsymbol{\theta}}$, implemented via one-step DDIM, to enforce cross-consistency throughout the ODE trajectory in both latent and pixel spaces. However, as reported in its paper, it suffers from unstable generation due to the excessive conditional guidance, leading to overexposure and artifacts. Furthermore, this approach results in a less stringent upper bound on the distillation error, expressed as $\mathcal{O}(\Delta_{t})(T-e)$, which constrains its generation quality.
Summary: The improper conditional guidance in CDS and GCS arises mainly from an imbalance between self-consistency and cross-consistency. Moreover, the reliance on $\boldsymbol{F_{\theta}}$ to enforce consistency across the entire trajectory further contributes to this issue.

### v2

@@ Abstract
Figure 1: Examples of 3D assets generated by SegmentDreamer. Our framework addresses the improper conditional guidance issues of recent Consistency Distillation (CD)-based methods and theoretically provides a significantly tighter upper bound on distillation error, enabling high-fidelity text-to-3D generation within a reasonable timeframe ($\sim$32 minutes with classifier-free guidance [9] and $\sim$38 minutes combined with Perp-Neg [1]) on a single A100 GPU through 3D Gaussian Splatting (3DGS). Project Page: https://zjhJOJO.github.io/segmentdreamer
footnotetext: *Corresponding author.
@@ 1 Introduction
Text-to-3D generation aims to create 3D assets that accurately align with text descriptions. It plays a vital role in various domains such as virtual reality, 3D gaming, and film production. One compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34] through Score Distillation Sampling (SDS) [31, 46], which leverages pre-trained text-to-image DMs as the 3D priors to guide the optimization of differentiable 3D representations, such as Neural Radiance Fields (NeRF) [29] and 3D Gaussian Splatting (3DGS) [17]. Subsequent research focuses on improving the performance of SDS in diversity [47, 11, 43], visual quality [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16], and multi-view consistency [5, 10, 37, 35, 12].
@@
Figure 2: Visual examples of existing CD-based methods: CDS [48], GCS [22], and our SegmentDreamer with a Classifier-Free Guidance (CFG) scale of 7.5. As shown, improper conditional guidance leads to suboptimal results. (a) Ineffective guidance produces semantically inconsistent details, while (b) excessive guidance results in overexposure and artifacts. Although (c) Brightness-Equalized Generation (BEG) [22] can mitigate overexposure, artifacts persist. In contrast, our SegmentDreamer enhances conditional generation, producing high-fidelity 3D assets.
@@
Recent methods [48, 22] have improved SDS and its variants by incorporating Consistency Distillation (CD) [40, 18] into their distillation loss. However, due to an imbalance between self-consistency and cross-consistency in their losses, these methods inherently suffer from improper conditional guidance when combined with Classifier-Free Guidance (CFG) for generation, ultimately leading to suboptimal outcomes (see Figure 2; a detailed explanation is provided in Section 4.1). Additionally, the distillation error upper bound of these methods remains large, limiting their generation quality.
@@
To address the aforementioned limitations, we present a novel framework – SegmentDreamer, which aims to fully unleash the potential of consistency models for high-fidelity text-to-3D generation. Specifically, by analyzing existing CD-based methods, we draw inspiration from segmented consistency models [33, 44] to reformulate SDS through the proposed Segmented Consistency Trajectory Distillation (SCTD), which effectively mitigates the imbalance between self-consistency and cross-consistency by explicitly defining their relationship. Moreover, unlike CD-based methods that directly enforce consistency across the entire Probability Flow Ordinary Differential Equation (PF-ODE) trajectory, SCTD partitions the PF-ODE trajectory into multiple sub-trajectories, ensuring consistency within each segment. This approach theoretically provides a significantly tighter upper bound on distillation error. In addition, we propose a distillation pipeline for a more swift and stable generation. We also introduce 3DGS as our 3D representation. Extensive experimental results demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality. Moreover, the proposed SCTD can be seamlessly applied to various 3D generative tasks (e.g., 3D avatar generation and 3D portrait generation), facilitating the creation of high-fidelity 3D assets.
@@ 2.1 Text-to-3D Generation
Creating 3D objects from text descriptions is challenging in the fields of computer vision and graphics. To achieve this goal, a series of end-to-end text-to-3D pipelines have been proposed [8, 7, 30, 13, 21, 49], but these methods require large-scale labeled 3D datasets for training. To reduce the reliance on 3D training data, a compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34]. A groundbreaking contribution in this domain is DreamFusion [31], which introduced the Score Distillation Sampling (SDS) loss to optimize parametric 3D representations. Follow-up works focus on refining and extending SDS in various aspects. For example, several studies [47, 11, 43] improve SDS in diversity through prompting modifications, while others [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16] enhance the fidelity of the generated 3D objects. Other advancements aim to achieve multi-view consistency [5, 10, 37, 35, 12] and faster generation [42, 52]. Recently, Consistency Distillation (CD) [40, 18] has received widespread attention as it enables high-quality generation through minimal sampling steps. Inspired by CD, Wu et al. [48] and Li et al. [22] incorporate CD into their distillation loss, which addresses some limitations brought by SDS. However, these CD-based methods inherently suffer from improper conditional guidance, leading to sub-optimal results. In this work, We aim to unleash CD’s potential in generating high-fidelity 3D assets and push text-to-3D synthesis boundaries.
@@ 2.2 Differentiable 3D Representations
Neural Radiance Field (NeRF) [29] aims to construct implicit 3D representations using a multi-layer perceptron (MLP), leveraging standard volumetric rendering [14] and alpha compositing techniques [32]. Subsequent research has explored adapting NeRF to various applications, including sparse-view reconstruction [45], anti-aliasing [2], and medical image super-resolution [4].
@@ 3 Preliminaries
Diffusion Model (DM) [39, 38, 15] consists of a diffusion process and a denoising process. The diffusion process can be modeled as the solution $\hat{\text{o}}$ stochastic differential equation (SDE) with the transition kernel $\mathbb{P}(\mathbf{z}_{t}|\mathbf{z}_{0}):=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$. By reversing the diffusion process, $\mathbf{z}_{0}$ can be estimated by solving a Probability Flow Ordinary Differentiable Equation (PF-ODE), which is defined by $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})]\mathrm{d}t$. A neural network ${\boldsymbol{\epsilon_{\phi}}}(\mathbf{z},t)$ can be trained to approximate the scaled score function $-\sigma_{t}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})$, offering an empirical formulation of the PF-ODE: $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}+\frac{g(t)^{2}}{2\sigma_{t}}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t)]\mathrm{d}t.$
Consistency Model [40, 27] aims to achieve few-step inference by distilling a pre-trained DM $\boldsymbol{\phi}$ into a student model $\boldsymbol{\phi}_{stu}$. Given a PF-ODE trajectory $\{\mathbf{z}_{t}\}_{t\in[0,T]}$ and a condition $\mathbf{y}$, a consistency model $\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}$ builds the mapping $\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}(\mathbf{z}_{t},t,\mathbf{y})=\mathbf{z}_{0}$ between any points along this trajectory and the endpoint $\mathbf{z}_{0}$ by minimizing the following Consistency Distillation (CD) loss:
$\mathcal{L}_{\text{CD}}=\mathbb{E}[\lambda(t)||\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}(\mathbf{z}_{t},t),\boldsymbol{f}_{\boldsymbol{\phi}_{stu}^{-}}(\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s)||_{2}^{2}],t>s,$ (1)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ for $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\boldsymbol{\phi}_{stu}^{-}$ is updated with the exponential moving average [40, 27]. $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ can be obtained by:
@@
The consistency function $\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}$ can be parameterized in various ways [40, 27], and the solver $\boldsymbol{\Phi}(\cdot,\cdot,\cdot,\cdot)$ is used to calculate the point at $s$, which can be implemented using DDIM [38] or DPM-solver [26].
Score Distillation Sampling (SDS) Loss. Given a 3D representation $\boldsymbol{\theta}$, one can render a view image $\mathbf{z}_{0}:=\boldsymbol{g}(\mathbf{c},\boldsymbol{\theta})$ with camera pose $\mathbf{c}$, SDS distills a 3D object from a 2D DM $\boldsymbol{\phi}$ by minimizing the following SDS loss w.r.t. $\boldsymbol{\theta}$:
@@
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$. Poole et al. [31] omit the Jacobian term $\frac{\partial\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\partial\mathbf{z}_{t}}$ in gradient calculation for speedup.
@@ 4 Our Method
In this section, we first review the mechanism of existing Consistency Distillation (CD)-based methods [48, 22] and reveal that they inherently suffer from improper conditional guidance when combined with Classifier-free Guidance (CFG). Subsequently, we address these limitations by reformulating SDS through the lens of the proposed Segmented Consistency Trajectory Distillation (SCTD). Moreover, building on this insight, we propose an efficient SCTD sampling method for high-fidelity text-to-3D generation. We further tailor a distillation pipeline for the proposed SCTD sampling to achieve more swift and stable optimization. Finally, we provide a theoretical proof for our distillation error upper bound. The overall framework of our SegmentDreamer is depicted in Fig. 3, with the algorithm detailed in Algorithm 1.
@@ 4.1 Reviewing CD-based Losses
The consistency function w.r.t. the 3D representation $\boldsymbol{\theta}$ is parameterized as $\boldsymbol{F_{\theta}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\mathbf{z}_{t}-\sigma_{t}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\alpha_{t}}$ [48, 22]. Before proceeding, we clarify two key concepts: self-consistency and cross-consistency. The former ensures that all points on the same ODE trajectory consistently map to their original points, as in consistency models [27, 40, 18]. The latter enforces alignment between unconditional and conditional ODE trajectories [22].
@@
[…]
@@
where $e$$<$$s$$<$$t$. GCS also employs a conditional guidance loss $\mathcal{L}_{\text{CG}}$ and a pixel-wise constraint $\mathcal{L}_{\text{CP}}$ to enforce cross-consistency (see Supp. A for details).
GCS ensures self-consistency within sub-trajectory, but it still suffers from improper conditional guidance due to two key issues: (1) $\boldsymbol{G_{\theta}}$ used to enforce cross-consistency is inherently flawed because $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$ lacks a target time-step, as detailed in App. B. (2) GCS employs $\boldsymbol{F}_{\boldsymbol{\theta}}$ (implemented via one-step DDIM) to enforce cross-consistency throughout the entire ODE trajectory in both latent and pixel spaces. However, as reported in its paper, it suffers from unstable generation due to the excessive conditional guidance, leading to overexposure and artifacts (more experimental analysis is provided in App. Fig. 1). Furthermore, this approach results in a less stringent upper bound on the distillation error, expressed as $\mathcal{O}(\Delta_{t})(T-e)$, which constrains its generation quality.
Table 1: Differences between CDS, GCS, and SCTD.“traj.” denotes trajectory.
CDS [48] GCS [22] SCTD (Ours)
Self-consistency ✓(whole traj.) ✓ (sub traj.) ✓ (sub traj.)
[…]
Error bound $\mathcal{O}(\Delta_{t})T$ $\mathcal{O}(\Delta_{t})(T\!-\!e)$ $\mathcal{O}(\Delta_{t})(s_{m+1}\!-\!s_{m})$
Summary: The improper conditional guidance in CDS and GCS stems mainly from an imbalance between self-consistency and cross-consistency. Additionally, their reliance on $\boldsymbol{F_{\theta}}$ to enforce consistency across the entire trajectory exacerbates this issue. Tab. 1 clarifies the differences between CDS, GCS, and ours.